OpenStudy (anonymous):

Runner A starts marathon going 5.6mph. Runner B starts a half hour later going 6.54mph. Marathon course is 26.2 mile long. which runner will finish first?

OpenStudy (kropot72):

Assume that timing begins for both runners when runner A starts. Time for runner A to finish = 26.2/5.6 = (you can calculate ) hrs Time for runner B to finish = (26.2/6.54) + 0.5 = (you can calculate ) hrs Obviously the runner with the shorter time will finish first.